ADNAN MUHAMMED

Grade 12,

The Referactive Index of glass and water with respect to air are 3/2 and 4/3 respectively. If speed of light in glass is 2*10 8 m/s, find the speed of light in water.

The Referactive Index of glass and water with respect to air are 3/2 and 4/3 respectively. If speed of light in glass is 2*108 m/s, find the speed of light in water.

Grade:10

1 Answers

Anwesha Barman
8 years ago
Absolute Refractive index of a  =  Speed of light in b 
Absolute Refractive index of b  =  Speed of light in a
Refractive indices in respect to air = Absolute refrative indices(n).
Therefore, n/ nb =  v/ va
 Or, (3/2) / (4/3) = vwater / 2 x 108.
Or, 9/8 x 2 x 108 = vwater.
Therefore, vwater = 9/4 x 108 m/s 2.25 x 10 m/s.